A population of values has a normal distribution with μ=206.3 and σ=51.5. You intend to draw a random sample of size n=202.
Find P64, which is the score separating the bottom 64% scores from the top 36% scores.
P64 (for single values) =
Find P64, which is the mean separating the bottom 64% means from the top 36% means.
P64 (for sample means) =
Enter your answers as numbers accurate to 1 decimal place.

Answer by Boreal(15235)

The z-value is +0.36 for the 64th percentile.
z=(x-mean)/sd or (x-mean)/sigma/sqrt(n)
0.36=(x-206.3)/51.5/sqrt(202)
0.36=(x-206.3)*sqrt(202)/51.5
18.54=(x-206.3)*sqrt(202)
1.304=x-206.3
x=207.6
check sd of sample distribution is roughly 51.5/14 or about 3.6
a third of 3.6 is about 1.2
it should be 1.2 above 206.3 and it is 1.3